This is a 10 part question in accounting, Thanks!

McCoys Fish House purchases a tract of land and an existing building for $980,000. The company plans to remove the old buildOrion Flour Mills purchased a new machine and made the following expenditures: Purchase price Sales tax Shipment of machine IRequired information The following information applies to the questions displayed below] Super Saver Grocerles purchased stor2. Double-declining-balance. Depreciation expense3. Activity-based Depreciation expense1. Straight-line. Year Annual Depreciation3. Activity-based Year Annual DepreciationTogos Sandwiches acquired equipment on April 1, 2021, for $15.500. The company estimates a residual value of $1.500 and a fiTasty Subs acquired a delivery truck on October 1, 2021, for $15,900. The company estimates a residual value of $2,100 and aTasty Subs acquired a delivery truck on October 1, 2021, for $16.500. The company estimates a residual value of $1.500 and a
Answer #1
1 Cost of the land:
$
Purchase price 980000
Title insurance 2800
Property taxes-Back taxes 8800
Cost of removing the building 49000
Sale of salvaged materials -4600
Leveling cost of land 10800
Total 1046800
2 Transaction General Journal Debit Credit
1 Equipment (Note:1) 77520
Prepaid insurance 640
Cash (940+640+1880) 3460
Accounts payable (69000+5700) 74700
(To record expenditure for the new machine)
Note:1 Cost of machinery:
$
Purchase price 69000
Sales tax 5700
Shipment of machine 940
Installation 1880
Total 77520
3
1. Straight-line:
Depreciation expense=(Cost-Salvage value)/Useful life=(28000-3000)/10=$ 2500
2. Double-declining balance:
Depreciation rate=2*(1/useful life)=2*(1/10)=2*0.10=0.20=20%
Depreciation for the first year=Cost*Depreciation rate=28000*20%=$ 5600
3. Activity-based:
Depreciation expense per hour=(Cost-salvage value)/Expected hours over the life=(28000-3000)/10000=$ 2.5 per hour
Depreciation for the first year=Depreciation expense per hour*Actual hours used=2.5*1680=$ 4200
4 Depreciation expense=(Cost-Salvage value)/Useful life=(15500-1500)/5=$ 2800
2021 2022
Depreciation expense 2800 2800
5 Depreciation expense=(Cost-Salvage value)/Useful life=(15900-2100)/6=$ 2300
2021 2022
Depreciation expense 2300 2300
6 Depreciation expense per miles=(Cost-salvage value)/Expected miles over the life=(16500-1500)/100000=$ 0.15 per hour
Depreciation for 2021=Depreciation expense per miles*Actual miles=0.15*4000=$ 600
Depreciation for 2022=Depreciation expense per miles*Actual miles=0.15*17000=$ 2550
2021 2022
Depreciation expense 600 2550
I appreciate your ratings